Weekly LIPAc highlights (23-27/November)

Injector conditioning

The Injector operation continued this week: the duty cycle was increased up to 30% with a very stable beam and 100mA of total extracted current from deuterium plasma. The vault will now be predominantly closed to continue the conditioning of the Injector in preparation of the facility inspection. The target performances for the facility inspection (planned on December 16) is 50% duty cycle at 100 mA of total extracted current.

Screenshot of the injector oscilloscope on 25 November:

  • Duty cycle: 31.2% (50ms/160ms)
  • Red: Total extracted current (Iext)
  • Yellow: Faraday cup signal (I_FC)
  • Blue: Current of repeller PS (I_repeller)
  • Green: Voltage of repeller PS (V_repeller, -4.5 kV)

RadioFrequency (RF) System

The RF team concluded several important activities this week:

  • The issue on the High Voltage Power supply 3A (HVPS3A) was solved with continuous remote support from JEMA.
  • The upgrade of the PSYS (Protection system) interlock was completed and installed on one RF module. The modification of the threshold levels of PSYS VSWR (Voltage Standing Wave Ratio) card were completed on the same module.
  • The anode box upgrade to decrease the overheating at the circuit board was successfully implemented in one module.

The plan is to start next week the entire RF system to validate the upgrades implemented in a part of the RF chains and to train QST personnel on the entire start-up process. Once the upgrades validated, the RF system will be stopped and the upgrades extended to all the modules.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc.